An Electron enters perpendicularly into a uniform magnetic field which has a flux density of 0.12T. This results in a magnetic force of 9.6x10-12N on the electron. Calculate the speed of the electron as it enters the magnetic field. [ e=1.6x10-19 C]
1
Expert's answer
2020-05-18T10:10:46-0400

"Force=q(v\times B )"


"9.6\times10^{-12}=1.6\times10^{-19}\times v\times0.12"


"v=5 \times 10^8m/s"